Let 𝑦1 and 𝑦2 be linearly independent solutions of the differential equation 𝑦 β€²β€² + 𝑝(π‘₯)𝑦 β€² + π‘ž(π‘₯)𝑦 = 0, where functions 𝑝 and π‘ž are continuous on some interval 𝐼. (i) Prove that π‘Š(𝑦1, 𝑦2 )(π‘₯) = 𝐢𝑒 βˆ’ ∫ 𝑝(π‘₯)𝑑π‘₯ , where π‘Š is the Wronskian and 𝐢 ∈ ℝ is an arbitrary constant.


1
Expert's answer
2021-11-15T07:43:25-0500

"y''+p(x)y''+q(x)y=0"

Since y1(x) and y2(x) are linear independent solutions of the differential equation above, we get:

"y_1''+p(x)y_1''+q(x)y_1=0" .........(1)

"y_2''+p(x)y_2''+q(x)y_2=0" ..........(2)

Now from (1) we get,

"y_1''=-(p(x)y_1'+y_1)" .......(3)

From (2) we get,

"y_2''=-(p(x)y_2'+y_2)" .........(4)

Now we know that the Wronskian

"W(y_1,y_2)=\begin{vmatrix}
 y_1 & y_2 \\
 y_1' & y_2'
\end{vmatrix}=y_1y_2'-y_2y_1'"

"W=y_1y_2'-y_2y_1'"

"W'=(y_1y_2'-y_2y_1')'"

"W'=(y_1y_2')'-(y_2y_1')'"

"W'=(y_1 \cdot y_2''+y_2'\cdot y_1')-(y_2\cdot y_1''+y_1'\cdot y_2')"

Since "(uv)'=u\cdot v'+v\cdot u'"

"W'=y_1 \cdot y_2''+y_2'\cdot y_1'-y_2\cdot y_1''-y_1'\cdot y_2'"

"W'=y_1\cdot[-p(x)y_2'+y_2)]-y_2\cdot[-(p(x)y_1'+y_1)]"

Values from (1)and (2)

"W'=-p(x)y_1y_2'-y_1y_2+p(x)y_2y_1'+y_1y_2"

"W'=-p(x)y_1y_2'+p(x)y_2y_1'"

"W'=-p(x)[y_1y_2'-y_2y_1']"

"W'=-p(x)W"

This is an equation of the form z'=-pz whose solution is in the form "z=ce^{-\int pdx}"

From "W'=-pW" we get

"W(y_1,y_2)=ce^{-\int Pdx}"